# Closed mean curvature flows with asymptotically conical singularities

### v2

@@ 5. Barrier functions
We look at the lower order terms of $\mathcal{Q}_{2}$ in (2.10). Using (2.9), if we denote $C=C(F,\nabla F,\nabla^{2}F,h,\nabla h)$ to be any tensor depending on $F,\nabla F,\nabla^{2}F,h,\nabla h$, we get
@@
Using (5.1), given $|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},$ we have
@@
$u=\begin{cases}e^{(-k+1)(\tau+\tau_{0})}\left(Df^{k}-Bf^{k-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}\\ e^{\tau+\tau_{0}}\left(a-Bf^{-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{\tau+\tau_{0}}<f(x)<\Gamma e^{\tau+\tau_{0}}\right\}\end{cases}.$
@@
$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
@@
$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-a\left(C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)-C_{0}\sqrt{a}\geq w>0,\text{ and}$
@@
$|h|_{C_{1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
@@
$\|h\|_{C^{1}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
$h(0)$ $=\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\text{ for }|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}$
$\|h(\tau)\|_{L_{W}^{2}(\Sigma)}$ $\leq\mu e^{\lambda_{*}(\tau+\tau_{0})}\text{ for all }\tau\in[0,\tau_{1}],$
$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$$\subseteq\left\{\frac{\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}}{2}\leq f\leq\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}\right\},$
@@
$|\nabla\mathcal{Q}_{1}(\cdot,\tau)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right).$
Then for any $\delta<\varepsilon$, $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma)$, $|\mathbf{p}|<C(n,\Sigma,F,\lambda_{*})$, $\mu<\mu_{0}(n,\Sigma,F,\lambda_{*})$, $\gamma<\gamma_{0}(n,\Sigma,F,\lambda_{*},\delta),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,\bar{p},\varepsilon,\delta)$, there exist $W=W(n,\Sigma,F,\lambda_{*})<\infty$ and $W^{\prime}=W^{\prime}(n,\Sigma,F)<\infty$ such that
$|h|+|\nabla h|+|\nabla^{2}h|$ $\leq W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}{(\tau+\tau_{0})}}\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}],\text{ and}$
$|h|+|\nabla h|+|\nabla^{2}h|$ $\leq W^{\prime}\delta\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}].$
@@ 6. Ważewski’s box argument
We are going to finish the proof of the main theorem in this section. For the concerned asymptotically conical shrinker $\Sigma,$ take $\tau_{0},\Gamma_{0}>0$ sufficiently large. From Section 3, we can construct $M=M_{\Sigma,\tau_{0},\Gamma_{0}},$ a doubling of a large compact set $\{f(x)\leq 2e^{\tau_{0}}\Gamma_{0}\}$ of $\Sigma$. Based on Proposition 2.3, we take an $L^{2}_{W}$-orthonormal eigenbasis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L$ on $\Sigma$ and the corresponding eigenvalues $\{\lambda_{i}\}_{i\in\mathbb{N}}$. Take $\lambda_{*}<0$ and $m$ such that $\lambda_{m+1}<\lambda_{*}<\lambda_{m}$.
@@
$F_{\mathbf{p}}(x,0)=F_{M}(x)+\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\cdot\textbf{n}_{F}(x).$
@@
where $F(x,t)$ corresponds to the family of immersions of mean curvature flows of the shrinker $\Sigma$. From the discussion in Section 2.2 we can take $F_{0}=\hat{F}_{\mathbf{p}}$. Then we get
@@
(6.1) $\left|d\hat{F}_{\mathbf{p}}(\partial_{t}\varphi_{t}(x))\right|\leq\frac{\left|(\partial_{t}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{\left|\left\langle\textbf{n}_{F}(x),\textbf{n}_{\hat{F}_{\mathbf{p}}}(\varphi_{t}(x))\right\rangle\right|}.$
@@
(6.2) $\frac{\tilde{h}^{ab}\left(1-|\nabla h|^{2}\right)-\left(|\nabla h|^{2}+\tilde{h}^{cd}\partial_{c}h\,\partial_{d}h\right)g^{ab}}{\left(1+\left\langle\textbf{n}_{F},\widetilde{\textbf{n}}\right\rangle\right)^{2}}\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}.$
@@
$\|h_{u}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{u}e^{\lambda_{*}(\tau+\tau_{0})},\text{ and}$
$\|h_{s}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{s}e^{\lambda_{*}{(\tau+\tau_{0})}}.$
@@
$|\nabla^{2}h|_{\bar{g}}$ $\leq\varepsilon_{2}.$
Our main goal here is to prove that for $\gamma_{0},|\mathbf{p}|<<1$ and $\Gamma_{0},\tau_{0}>>1$, $h_{\mathbf{p}}$ can only leave $\mathcal{B}$ from the unstable side.
@@
Given $\gamma_{0},|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, for $h_{\mathbf{p}}$ constructed above, assume $h_{\mathbf{p}}\notin\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ If we denote the largest existence time for $F_{\mathbf{p}}(x,\cdot)$ by $T(\mathbf{p})$ and take the largest time $t(\mathbf{p})$ such that $h_{\mathbf{p}}\in\mathcal{B}=\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},t(\mathbf{p})]$, then we have
$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
@@
From the standard ordinary differential equation and parabolic theories, we have that for $|\mathbf{p}|<<1$, we can get $t(\mathbf{p})>0$. From the interior estimate and the pseudolocality in [EH, INS], on $M\setminus\left\{f\leq\frac{\Gamma_{0}e^{\tau_{0}}}{2}\right\}$, $F_{\mathbf{p}}(x,t)$ satisfies $|\nabla^{m}A|\leq C_{m}(\frac{1}{\Gamma_{0}e^{\tau_{0}}})^{\frac{m+1}{2}}$ for $t\in[0,2]$. Thus, if $h_{\mathbf{p}}\in\mathcal{B}$, we have
@@
Hence, if we denote $\tau=-\log(1-t)$,
$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$ $=\text{supp}\left({\sqrt{1-t}}\left\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(\cdot),t),\textbf{n}_{F_{1}}(\cdot)\right\rangle\right)\subseteq\left\{\frac{\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}}{2}\leq f\leq\Gamma_{0}e^{\frac{(\tau+\tau_{0})}{2}}\right\},$
@@
$\|h\|_{C^{2}(\Sigma\times[0,\tau_{1}])}\leq W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}(\tau+\tau_{0})}.$
@@
$|\mathcal{Q}_{2}|\leq C\left(|h|+|\nabla h|+|\nabla^{2}h|\right)^{2}\leq C\min\left\{W\left(\frac{|F|^{2}}{4}\right)^{|\lambda_{*}|}e^{\lambda_{*}(\tau+\tau_{0})},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\right\}.$
@@
(6.3) $\|\mathcal{Q}_{1}\|_{L^{2}_{W}}+\|\mathcal{Q}_{2}\|_{L^{2}_{W}}\leq CW^{2}e^{2\lambda_{*}(\tau+\tau_{0})}\leq Ce^{2\lambda_{*}(\tau+\tau_{0})}.$
From the definition of $h_{\mathbf{p}}$, the polynomial decay of eigenfunctions (Theorem 2.4), and the exponential decay of the weight, for $|\mathbf{p}|\leq\bar{p}e^{\lambda_{*}\tau_{0}}$, we have
@@
(6.5) $\frac{d}{d\tau}\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}$ $\geq(\lambda_{m}-\lambda_{*})\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}-Ce^{\lambda_{*}\tau_{0}},\text{ and}$
$\frac{d}{d\tau}\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},s}\|_{L^{2}_{W}}$ $\leq(\lambda_{m+1}-\lambda_{*})\|e^{-\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}+Ce^{\lambda_{*}\tau_{0}}.$
@@
$\|h_{\mathbf{p},s}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}<\mu_{s}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))},$
@@
$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
@@
$\|h_{\mathbf{p},u}(\cdot,\tau)\|_{L^{2}_{W}}>\mu_{u}e^{\lambda_{*}(\tau+\tau_{0})}.$
@@
Given $\gamma_{0},$ $\bar{p}e^{\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$ such that
@@
As in the Ricci flow case, we define $\mathcal{F}:\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}$ by
@@
$\mathcal{F}|_{\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}}\colon\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$
is homotopic to the identity through $\mathcal{F}_{s}:=s\mathcal{F}+(1-s){\rm Id}\colon\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$.
On the other hand, if
@@
for any $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}},$ then Proposition 6 implies that $|\mathcal{F}(\mathbf{p})|>0$ for $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$. This shows that, in particular,
$\mathcal{F}|_{\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}}:\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}\setminus\overline{B_{2\mu_{u}e^{\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}\setminus\{0\}$
@@
Hence, we derive a contraction. This proves that there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{\lambda_{*}\tau_{0}}}}$ such that
@@
Given an asymptotically conical shrinker $\Sigma^{n}\subseteq\mathbb{R}^{n+1}$, there exists a closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
@@
For any $\mathbf{p},$ if we blow up the mean curvature flow starting from the corresponding perturbation at $({\bf 0},1)$, the family of rescaled hypersurfaces is the graph of $h_{\mathbf{p}}$ over $\Sigma$. From Proposition 6, there exists a vector $\mathbf{p}$ such that $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ Based on Theorem 5 and Proposition 4, $h_{\mathbf{p}}$ converges to 00 in $C^{\infty}_{loc}(\Sigma)$. The global $C^{2}$ bound of $h_{\mathbf{p}}$ guarantees that the singularity is of type I. This finishes the proof. ∎

### v3

@@ 5. Barrier functions
We look at the lower order terms of $\mathcal{Q}_{2}$ in (2.9). Using (2.8), if we denote $C=C(F,\nabla F,\nabla^{2}F,h,\nabla h)$ to be any tensor depending on $F,\nabla F,\nabla^{2}F,h,$ and $\nabla h$, we get
@@
Using (5.1), given $\|h\|_{C^{1,\alpha}_{\rm hom,-1}}\leq\varepsilon<\varepsilon_{n},$ we have
@@
$u=\begin{cases}e^{(-k+1)(\tau+\tau_{0})}\left(Df^{k}-Bf^{k-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\Gamma<f(x)<\gamma_{1}e^{\tau+\tau_{0}}\right\}\\ e^{\tau+\tau_{0}}\left(a-Ef^{-1}\right)&\text{on }\left\{(x,\tau)\in\Sigma\times[0,\infty):\gamma_{2}e^{\tau+\tau_{0}}<f(x)<\Gamma e^{\tau+\tau_{0}}\right\}\end{cases}$
for some positive numbers $a,B,D,$ and $E.$
@@
$\|h\|_{C^{1,\alpha}_{\rm hom,-1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m\left(m+\frac{n}{2}\right)+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0,$
@@
$|h|_{C^{\rm hom}_{1;1}}\leq\varepsilon<\varepsilon_{n},\quad B-a\left(C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)-C_{0}\sqrt{a}\geq w>0,\text{ and}$
@@
$\|h\|_{C^{1,\alpha}_{\rm hom,-1}}\leq\varepsilon<\varepsilon_{n},\quad B-D\left(m(m+\frac{n}{2})+C_{n}\sup_{\Sigma}\left(f|A|^{2}+f|\nabla A|\right)\right)\geq w>0$
@@
$\|h\|_{C^{1,\alpha}_{\rm hom,-1}(\Sigma\times[0,\tau_{1}])}$ $<\varepsilon,$
$h(0)$ $=\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\text{ for }|\mathbf{p}|\leq\bar{p}e^{-\lambda_{*}\tau_{0}},$
$\|h(\tau)\|_{L_{W}^{2}(\Sigma)}$ $\leq\mu e^{-\lambda_{*}(\tau+\tau_{0})}\text{ for all }\tau\in[0,\tau_{1}],$
$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$$\subseteq\left\{\frac{\Gamma_{0}e^{(\tau+\tau_{0})}}{4}\leq f\leq\Gamma_{0}e^{(\tau+\tau_{0})}\right\},$
@@
$|\nabla\mathcal{Q}_{1}(\cdot,\tau)|$ $\leq C\left(\frac{1}{\Gamma_{0}e^{\tau+\tau_{0}}}\right)$
for some $\lambda_{*}>0.$ Then for any $\delta<\varepsilon$, $\Gamma>\Gamma_{0}(n,\Sigma,F)$, $\varepsilon<\varepsilon_{0}(n,\Sigma,F,\Gamma)$, $|\mathbf{p}|<C(n,\Sigma,F,\lambda_{*})$, $\mu<\mu_{0}(n,\Sigma,F,\lambda_{*})$, $\gamma<\gamma_{0}(n,\Sigma,F,\lambda_{*},\delta),$ and $\tau_{0}>C(n,\Sigma,F,\lambda_{*},\Gamma,\bar{p},\varepsilon,\delta)$, there exist $W=W(n,\Sigma,F,\lambda_{*})<\infty$ and $W^{\prime}=W^{\prime}(n,\Sigma,F)<\infty$ such that
$|\nabla^{l}h|$ $\leq W\tilde{r}^{2\lambda_{*}+1-l}e^{-\lambda_{*}{(\tau+\tau_{0})}},\,\,l\in\{0,1,2\}\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}],\text{ and}$
$|\nabla^{l}h|$ $\leq W^{\prime}\delta\tilde{r}^{1-l},\,\,l\in\{0,1,2\}\text{ for }(x,\tau)\in\Sigma\times[0,\tau_{1}].$
@@ 6. Ważewski’s box argument
We are going to finish the proof of the main theorem in this section. For the concerned asymptotically conical shrinker $\Sigma,$ take $\tau_{0},\Gamma_{0}>0$ sufficiently large. From Section 3, we can construct $M=M_{\Sigma,\tau_{0},\Gamma_{0}},$ a doubling of a large compact set $\{f(x)\leq 2e^{\tau_{0}}\Gamma_{0}\}$ of $\Sigma$. Based on Proposition 2.3, we take an $L^{2}_{W}$-orthonormal eigenbasis $\{h_{i}\}_{i\in\mathbb{N}}$ of $L$ on $\Sigma$ and the corresponding eigenvalues $\{\lambda_{i}\}_{i\in\mathbb{N}}$. Take $\lambda_{*}>0$ and $m$ such that $\lambda_{m}<\lambda_{*}<\lambda_{m+1}$.
@@
$F_{\mathbf{p}}(x,0)=F_{M}(x)+h_{\mathbf{p}}(x)\textbf{n}_{F}(x):=F_{M}(x)+\eta\left(\frac{1}{\gamma_{0}e^{\tau_{0}}}f(x)\right)\sum_{i=1}^{m}p_{i}f_{i}\cdot\textbf{n}_{F}(x).$
@@
where $F(x,t)$ corresponds to the family of immersions of mean curvature flows of the shrinker $\Sigma$. From the discussion in Section 2.2, we can take $F_{0}=\hat{F}_{\mathbf{p}}$. Then we get
@@
$\left|d\hat{F}_{\mathbf{p}}(\partial_{t}\varphi_{t}(x))\right|\leq\frac{\left|(\partial_{t}\hat{F}_{\mathbf{p}})(\varphi_{t}(x),t)+\frac{1}{2\sqrt{1-t}}F(x)\right|}{\left|\left\langle\textbf{n}_{F}(x),\textbf{n}_{\hat{F}_{\mathbf{p}}}(\varphi_{t}(x))\right\rangle\right|}.$
@@
$\frac{\tilde{h}^{ab}\left(1-|\nabla h|^{2}\right)-\left(|\nabla h|^{2}+\tilde{h}^{cd}\partial_{c}h\,\partial_{d}h\right)g^{ab}}{\left(1+\left\langle\textbf{n}_{F},\widetilde{\textbf{n}}\right\rangle\right)^{2}}\partial_{a}\partial_{b}h\cdot\textbf{n}_{F}.$
@@
$\|h_{u}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{u}e^{-\lambda_{*}(\tau+\tau_{0})},\text{ and}$
$\|h_{s}(\cdot,\tau)\|_{L^{2}_{W}}$ $\leq\mu_{s}e^{-\lambda_{*}{(\tau+\tau_{0})}}.$
@@
$|\nabla^{2}h|_{\bar{g}}$ $\leq\varepsilon_{2}\tilde{r}^{-1}.$
where $\tilde{r}$ is a fixed smooth function such that $\tilde{r}\geq 1$ and $\tilde{r}=|x|$ for $|x|\geq 2$. Our main goal here is to prove that for $\gamma_{0},|\mathbf{p}|<<1$ and $\Gamma_{0},\tau_{0}>>1$, $h_{\mathbf{p}}$ can only leave $\mathcal{B}$ from the unstable side.
@@
Given $\gamma_{0},|\mathbf{p}|\leq\bar{p}e^{-\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, for $h_{\mathbf{p}}$ constructed above, assume $h_{\mathbf{p}}\notin\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ If we denote the largest existence time for $F_{\mathbf{p}}(x,\cdot)$ by $T(\mathbf{p})$ and take the largest time $t(\mathbf{p})$ such that $h_{\mathbf{p}}\in\mathcal{B}=\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},t(\mathbf{p})]$, then we have
$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{-\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
@@
From the standard ordinary differential equation and parabolic theories, for $|\mathbf{p}|<<1$, we can get $t(\mathbf{p})>0$. From the interior estimate and the pseudolocality in [EH, INS], on $M\setminus\left\{f\leq\frac{\Gamma_{0}e^{\tau_{0}}}{2}\right\}$, $F_{\mathbf{p}}(x,t)$ satisfies $|\nabla^{m}A|\leq C_{m}(\frac{1}{\Gamma_{0}e^{\tau_{0}}})^{\frac{m+1}{2}}$ for $t\in[0,2]$. Thus, if $h_{\mathbf{p}}\in\mathcal{B}$, we have
@@
Hence, if we denote $\tau=-\log(1-t)$, we have
$\text{supp}(\mathcal{Q}_{1}(\cdot,\tau))$ $=\text{supp}\left({\sqrt{1-t}}\left\langle(\partial_{t}\hat{F}_{\mathbf{p}}-\Delta^{g_{0}}\hat{F}_{\mathbf{p}})(\varphi_{t}(\cdot),t),\textbf{n}_{F_{1}}(\cdot)\right\rangle\right)\subseteq\left\{\frac{\Gamma_{0}e^{(\tau+\tau_{0})}}{2}\leq f\leq\Gamma_{0}e^{(\tau+\tau_{0})}\right\},$
@@
$\|h\|_{C^{2}(\Sigma\times[0,\tau_{1}])}$ $\leq W\left(\frac{|F|^{2}}{4}\right)^{\lambda_{*}}e^{-\lambda_{*}(\tau+\tau_{0})}.$
@@
$|\mathcal{Q}_{2}|\leq C\left(|h|+|\nabla h|+|\nabla^{2}h|\right)^{2}\leq C\min\left\{W\left(\frac{|F|^{2}}{4}\right)^{\lambda_{*}}e^{-\lambda_{*}(\tau+\tau_{0})},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\right\}.$
@@
(6.1) $\|\mathcal{Q}_{1}\|_{L^{2}_{W}}+\|\mathcal{Q}_{2}\|_{L^{2}_{W}}\leq CW^{2}e^{-2\lambda_{*}(\tau+\tau_{0})}\leq Ce^{-2\lambda_{*}(\tau+\tau_{0})}.$
From the definition of $h_{\mathbf{p}}$, the polynomial decay of eigenfunctions (Theorem 2.4), and the exponential decay of the weight, for $|\mathbf{p}|\leq\bar{p}e^{-\lambda_{*}\tau_{0}}$, we have
@@
(6.3) $\frac{d}{d\tau}\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}$ $\geq(\lambda_{*}-\lambda_{m})\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}-Ce^{-\lambda_{*}\tau_{0}},\text{ and}$
$\frac{d}{d\tau}\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},s}\|_{L^{2}_{W}}$ $\leq(\lambda_{*}-\lambda_{m+1})\|e^{\lambda_{*}(\tau+\tau_{0})}h_{\mathbf{p},u}\|_{L^{2}_{W}}+Ce^{-\lambda_{*}\tau_{0}}.$
@@
$\|h_{\mathbf{p},s}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}<\mu_{s}e^{-\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))},$
@@
$\|h_{\mathbf{p},u}(\cdot,-\log(1-t(\mathbf{p})))\|_{L^{2}_{W}}=\mu_{u}e^{-\lambda_{*}(\tau_{0}-\log(1-t(\mathbf{p})))}.$
@@
$\|h_{\mathbf{p},u}(\cdot,\tau)\|_{L^{2}_{W}}>\mu_{u}e^{-\lambda_{*}(\tau+\tau_{0})}.$
@@
Given $\gamma_{0},$ $\bar{p}e^{-\lambda_{*}\tau_{0}}<<1$ and $\Gamma_{0},\tau_{0}>>1$, there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$ such that
@@
As in the Ricci flow case, we define $\mathcal{F}:\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\to\mathbb{R}^{m}$ by
@@
$\mathcal{F}|_{\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}}\colon\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}\to\mathbb{R}^{m}\setminus\{0\}$
is homotopic to the identity through $\mathcal{F}_{s}:=s\mathcal{F}+(1-s){\rm Id}\colon\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}\to\mathbb{R}^{m}\setminus\{0\}$.
We claim that this shows the Proposition. Otherwise, if the proposition is false, then
@@
for all $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$. Thus Proposition 6 implies that $|\mathcal{F}(\mathbf{p})|>0$ for all $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$. Note that $\mathcal{F}|_{\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}}$ is null-homotopic to a constant map by shrinking $\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$ to a point and compose it with $\mathcal{F}$. Here, we have that the whole homotopy lies in $\mathbb{R}^{m}\setminus\{0\}$ as $|\mathcal{F}(\mathbf{p})|>0$ for all $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$. This shows that, in particular,
$\mathcal{F}|_{\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}}\colon\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}\setminus B_{2\mu_{u}e^{-\lambda_{*}\tau_{0}}}\to\mathbb{R}^{m}\setminus\{0\}$
@@
Hence, we derive a contraction. This proves that there exists $\mathbf{p}\in\overline{B_{\bar{p}e^{-\lambda_{*}\tau_{0}}}}$ such that
@@
Given an asymptotically conical shrinker $\Sigma^{n}\subseteq\mathbb{R}^{n+1}$, there exists a closed embedded hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
@@
For any $\mathbf{p},$ if we blow up the mean curvature flow starting from the corresponding perturbation at $({\bf 0},1)$, the family of rescaled hypersurfaces is the graph of $h_{\mathbf{p}}$ over $\Sigma$. From Proposition 6, there exists a vector $\mathbf{p}$ such that $h_{\mathbf{p}}\in\mathcal{B}[\lambda_{*},\mu_{u},\mu_{s},\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},1].$ Based on Theorem 5 and Proposition 4, $h_{\mathbf{p}}$ converges to 00 in $C^{\infty}_{\rm loc}(\Sigma)$. The global $C^{2}$ bound of $h_{\mathbf{p}}$ guarantees that the singularity is of type I. This finishes the proof. ∎